# Find Any Part in 3D

### v1

@@ Abstract
We study open-world part segmentation in 3D: segmenting any part in any object based on any text query. Prior methods are limited in object categories or part vocabularies. Recent advances in AI have demonstrated effective open-world recognition capabilities in 2D. Inspired by this progress, we propose an open-world, direct-prediction model for 3D part segmentation that can be applied zero-shot to any object. Our approach, called Find3D, trains a general-category point embedding model on large-scale 3D assets from the internet without any human annotation. It combines a data engine, powered by foundation models for annotating data, with a contrastive training method. We achieve strong performance and generalization across multiple datasets, with up to a 3$\times$ improvement in mIoU over the next best method. Our model is 6$\times$ to over 300$\times$ faster than existing baselines. To encourage research in general-category open-world 3D part segmentation, we also release a benchmark for general objects and parts. Project website: https://ziqi-ma.github.io/find3dsite/
Figure 1: Find3D is an open-world part segmentation model that can segment any part of any object with any text query. Our method consists of a Data Engine powered by 2D foundation models – SAM & Gemini – that automatically annotates 3D assets from the web. Using the labeled data, Find3D trains a transformer-based point cloud model with a contrastive-based training recipe. Our method can segment diverse 3D objects with a rich set of part queries, such as the easel, as well as in-the-wild 3D reconstructions obtained via image-to-3D methods, such as the push car captured with an iPhone and the capybara with a hat generated by DALL-E, shown at the bottom right.
@@ 1 Introduction
We study open-world part segmentation in 3D. By open-world, we mean that: 1) The input can be any object. 2) the model should handle any text query, enabling flexibility in granularity (e.g., “limbs” versus “arm”) and description type (e.g., “glove” versus “hand”). Recent advancements in foundation models show that single models can achieve an impressive understanding of general concepts in language and 2D images. These results beg the question: how far are we from a foundation model that understands the 3D world? To answer this question, we first ask: can we build a model that can find any part of any object in 3D? Such a capability is critical to robotic and VR/AR applications where grounding objects and their parts in 3D is necessary for precise manipulation, interaction, and spatial awareness. We study the point cloud representation, which is the most readily available representation in such applications.
Segmentation is a longstanding problem in computer vision. However, the problem has generally been formulated in a closed world, and at the whole object level (e.g., “car”). Some efforts tackle perception below the object level (e.g., “wheel of a car”), mainly driven by ShapeNet-Part [29], a closed-world benchmark of 16 object classes such as tables and chairs, with 6 or fewer predefined parts per object.
Recent vision and language foundation models have shown impressive zero-shot generalization capabilities in 2D image understanding  [18, 1]. These models are impressive as they operate in an open-world setting. Unfortunately, they do not apply to 3D representations. To build an effective open-world 3D model, 3D training data is needed for a diverse set of object types, shapes, and parts. However, it is not easy to acquire such data.
In this paper, we develop a data engine for automatically labeling online 3D assets, and train an open-world 3D part segmentation model. The data engine is powered by 2D vision and language foundation models and annotates 3D objects with diverse part labels in an automated fashion. Using the data collected by the data engine, we train a transformer-based 3D model that takes in a point cloud and predicts queryable semantic features for every point. The features are in the latent embedding space of a CLIP-like [18] model, so that they can be queried with any free-form text by calculating pointwise cosine similarities with the query embedding. We devise a contrastive training objective to handle part hierarchy and ambiguity. Fig. 1 shows the overall schematic of our approach, which we call Find3D.
Find3D, a model of $46.2M$ total parameters trained on $27K$ labeled objects, can find any part of any object. We highlight the following contributions:
  • We build a zero-shot, part-level, open-world, direct-prediction model for 3D point cloud semantic segmentation that works on general object categories and part queries. We achieve 3$\times$ the mIoU and 6$\times$ to over 300$\times$ the inference speed compared to existing methods.
@@
  • We release a benchmark for evaluating open-world part-level semantic segmentation for general object categories with a diversity of objects and no pose constraints.
Figure 2: The Data Engine. We render Objaverse assets into multiple views and pass each rendering to SAM with gridpoint prompts for segmentation. For each mask, we query Gemini for the corresponding part name, which gives us (mask, text) pairs. We embed the part name into the latent embedding space of a vision and language foundation model such as SigLIP. We back-project mask pixels to obtain the points associated with each label embedding, yielding (points, text embedding) pairs as visualized on the right side of the figure.
@@ 2 Related Work
Closed-world 3D segmentation. 3D segmentation has been studied primarily in a closed world and with a coarse granularity that cannot go below whole objects. In specific settings such as indoor scenes or self-driving, state-of-the-art models are starting to achieve better generalization by training on multiple datasets, such as Mask3D [20] and the PointTransformer series [32, 27, 26]. However, these models are still limited to their benchmark settings, and can only segment whole objects rather than parts. Part-level segmentation is less studied. Early efforts started with the ShapeNet-Part dataset [29] (16 object classes, $\leq 6$ parts per object). PartNet-E, a more recent part-level dataset combining PartNet-Mobility [28] and PartNet [14], introduces articulated objects but is still limited to only 45 categories. Due to the limited number of categories and shared orientations (e.g., chairs all facing right), state-of-the-art part-level models [17, 11] cannot generalize well. Our work tackles both the challenge of generalization and granularity – our model is designed to be part-level, and can segment any object part in an open-world setting.
3D aggregation methods based on 2D renderings. With the progress of vision language models in 2D image understanding, some works directly assemble these models to obtain an “aggregated” 3D understanding without training a 3D model. An exemplary aggregation method uses multiview renderings of 3D scenes or objects, obtains their features in 2D based on models like CLIP [18], SAM [8], or GLIP [9], and combines them in 3D based on projection geometry. On the whole object level, such methods include OpenMask3D [22]. On the part level, such methods include PointCLIP [31], PointCLIPV2 [35], PartSLIP [10] for point clouds, and PartSLIP++ [34] for meshes. Because these models take features only from 2D, they lack 3D geometry information and suffer from inconsistency across views. Furthermore, these methods are slow because they perform many inferences and the aggregation logic at test time. Our method works natively on 3D point clouds without an aggregation pipeline, and is significantly faster. Our method also uses 3D geometry information to achieve stronger performance and better robustness to pose changes.
Test-time optimization. Test-time optimization methods combine features from 2D models with a 3D representation, such as NeRF or Gaussian Splatting. At test time, these methods optimize the 3D representation with the 2D-sourced features attached. LERF [5], Distilled Feature Field [21], and Garfield [6] are based on radiance fields. Feature3DGS [33] is based on Gaussian splatting. These methods can provide good performance, but need to be optimized per scene (or per object), which can take several minutes. Besides being slow, the part-level capabilities have not been well-studied for such methods. Our method, being a feed-forward 3D model designed for part segmentation, provides much faster inference with better performance.
@@ 3 Method
Figure 3: Find3D: an open-world part segmentation model. Find3D takes in a point cloud, voxelizes and serializes the points via space-filling curves into a sequence. The sequence is passed through a transformer architecture which performs block attention, shuffling, pooling and unpooling, returning a pointwise feature that is in the embedding space of a vision and language foundation model, denoted by $\mathcal{T}$. These features can be queried with any free-form text. Find3D is trained with a contrastive objective. For each (points, text embedding) label from the data engine, we use the averaged feature of these points as the predicted embedding, and pair it with the text embedding to form a positive pair in the contrastive loss.
@@
Formally, given a point cloud $C=\{\mathbf{p_{i}},...{\mathbf{p_{n}}}\}$ with color and normals, for any point $\mathbf{p_{i}}=(x,y,z,n_{x},n_{y},n_{z},r,g,b)\in\mathbb{R}^{9}$, we want to find a semantic feature $f_{i}\in\mathbb{R}^{d}$ which belongs in the same latent embedding space as a CLIP-like model, e.g., SigLIP. These semantic features attached to the point cloud $C$ allow us to query the point cloud in 3D space in an open-world fashion – for any text $s$, we can get its SigLIP embedding $\mathcal{T}(s)$ and compute its cosine similarity with $f_{i}$, $cos(\mathcal{T}(s),f_{i})$.
To obtain such a model, Find3D consists of a data engine and a contrastive training recipe. The data engine automatically annotates online 3D assets, leveraging 2D foundation models. These annotations are used to train a Transformer-based [24] model for 3D point clouds. We use a contrastive learning objective to overcome the challenge of label ambiguity (e.g. one point associated with “snowball” and “body” from different views). Our data engine and contrastive training recipe enable scalable training on diverse objects and part labels.
@@ 3.1 Data Engine
Obtaining large-scale 3D annotations for generic object categories with human-in-the-loop pipelines is onerous. To this end, we develop a scalable data engine that leverages annotations from 2D foundation models and geometrically unprojects them to 3D.
As illustrated in  Fig. 2, Our data engine leverages SAM [8] and Gemini [19] to annotate 3D assets from Objaverse [3]. Since Objaverse assets do not have a fixed orientation, and Gemini provides higher-quality labels to objects seen in familiar orientations, we first prompt Gemini to select the best orientation based on 10 renderings (from different camera angles) of an object in each orientation. For the chosen orientation, we pass all renderings to SAM with grid point prompts. We discard masks that are too small (less than 350 pixels out of a 500x500 image), too large (greater than 20% of all pixels), or with low confidence from SAM. We overlay each mask on the original image and ask Gemini to name the shaded part. Prompts are detailed in the appendix. Masks from the same camera view with the same label are merged. This process generates labeled (mask, text) pairs. We map each mask to a set of points in the point cloud based on projection geometry. To make the point features queryable by language, we align point features to the language embedding space of a pretrained model, such as SigLIP. We embed the label texts and use the text features as supervision. Thus, our label (mask, text) pairs become (points, text embedding) pairs.
The data engine processes 36044 high-quality objects of Objaverse under LVIS categories selected by  [12, 13]. After filtering out objects with insufficient labels, we obtain 30K annotated objects from 761 categories with 1.5 million part annotations. Each part is annotated differently from different views, denoting various aspects of part, such as location (e.g., “bottom”), material (e.g., “snowball”), and function (e.g., “body”). Labels also have different levels of granularity. For example, in Fig. 2, one granularity is individual snowballs, and another granularity is the whole snowman. The diversity of our labels helps the model handle the inherent ambiguity in segmentation.
@@ 3.2 Open-World 3D Part Model
As shown in Fig. 3, Find3D takes in a point cloud and returns a pointwise semantic feature that can be queried with any free-form text via cosine similarity. For any number of user-specified text queries, Find3D first embeds them via the SigLIP text encoder. Find3D then computes a pairwise cosine similarity of each point and the text queries. To produce 3D segmentations, Find3D assigns each point to the text query with the highest cosine similarity, and assigns “no label” if all queries yield negative similarity scores.
3.2.1 Architecture
Find3D adopts a point transformer architecture that treats point clouds as sequences. We build upon the PointTransformer3 (PT3) model [26]. Our design is illustrated in Fig. 3. First, the point clouds are converted into sequences. To this end, the normalized point clouds are sampled with a voxel size of 0.02, so that there is at most one point per voxel. The voxels can be serialized in different methods, including Z-Ordering, Trans-Z-Ordering, Hilbert, and Trans-Hilbert [15, 4, 26]. These different serialization schemes are used in different layers in our architecture, following PT3. The serialized point cloud forms a sequence of length $N$, where $N$ is the number of points. Each point carries its XYZ, RGB, and normal information. These values are embedded with a linear layer into point embedding, and a conditional positional encoding [2] is added to the embedding. The embedded sequence is then passed to the transformer, which comprises an encoder and a decoder.
Since the point sequence can be arbitrarily long and the quadratic complexity of attention can grow prohibitively expensive, we chunk the sequence into 1024-size blocks as in the PT3 architecture. Attention is only performed within each block. To achieve the large receptive field of transformers, shuffling is applied per layer. Each shuffling operation maps the current serialization scheme to the (different) serialization scheme used by the next layer. In the encoder, pooling is applied per layer. Although pooling is dependent on the serialization scheme, the goal is to pool points that are neighbors in the 3D space, since all our serialization schemes preserve locality (to the extent possible). As a result, the sequence becomes shorter, and the receptive field becomes larger, even though block attention has a fixed size. In the decoder, this process is reversed – unpooling happens by back-tracing to pooled points in the encoding stage. Block attention happens similarly to the encoder, while shuffling is also back-traced. Eventually, we recover a sequence of the same length as the input (i.e., number of points), but of a different latent dimension. Since we want to align the point feature into the latent embedding space of SigLIP, we append a lightweight 4-layer MLP to the last layer of the transformer. This returns a 768-dimension feature per point. Our model contains 46.2 million parameters.
During training, the contrastive loss is only applied to points that are kept during voxel sampling. At inference time, the model only predicts features for the points that are kept during voxel sampling. For the points that are dropped, we assign the feature of the nearest neighbor which is kept during voxel sampling.
3.2.2 Training
@@
The data engine provides diverse part labels. However, it is difficult to define a direct pointwise loss because: 1) The same point can have multiple labels that denote various aspects of a part such as location, material, and function. Some labels may also be incorrect. 2) Many points are unlabeled - as shown in Fig. 3 (right), each mask only labels points visible from one camera view.
@@
Due to these challenges, we resort to a contrastive loss that does not rely on per-point labeling. As illustrated on the right side of Fig. 3, we define the pairing as follows: for each label, the ground truth is the SigLIP embedding of the text. The predicted value is taken as the average feature of all points belonging to the label. This pooling can also be regarded as a way to “denoise” the labels – while an individual point might be affected by conflicting or incorrect labels, it is unlikely that all points are subjected to the same conflicting or incorrect labels.
@@
Our data engine provides (points, text embedding) labels, which we denote as $(C_{i},\mathcal{T}(\text{label}_{i}))$ where $C_{i}$ is a subset of the point cloud that this label applies to, and $\mathcal{T}(\text{label}_{i})$ is the label embedding. We denote the pooled feature from the labeled points as $f(C_{i})$, where $f$ is our model. We define the contrastive loss as follows:
$l_{i}=-\log\frac{\exp(f(C_{i})\cdot\mathcal{T}(\text{label}_{i}))}{\sum_{j=1}^{|\mathcal{B}|}\exp(f(C_{i})\cdot\mathcal{T}(\text{label}_{j}))}$ (1)
@@
where $\mathcal{B}$ is the indices of all labels of all objects in a batch.
Because each label is unprojected from a single view, they only include points visible from that view and cover a part only partially. Although this illustration only shows one object, during training we include $64$ objects in the batch, corresponding to $\sim 3000$ positive pairs per batch.
@@
To achieve generalization, in addition to training on diverse data provided by the data engine, we also apply data augmentations, including full random rotation (implemented as sequential random rotation along all three axes), scaling, flipping, jittering, chromatic auto contrast, chromatic translation, and chromatic jitter. These augmentations help avoid over-reliance on object poses and color, and nudges the model to take up 3D geometric cues. We perform a 90:10 train-validation split on the 27552 objects provided by the data engine, and train with the Adam optimizer [7] with a cosine annealing learning rate schedule, starting at 0.0003 and ending at 0.00005 over 80 epochs.

### v2

@@ Abstract
Why don’t we have foundation models in 3D yet? A key limitation is data scarcity. For 3D object part segmentation, existing datasets are small in size and lack diversity. We show that it is possible to break this data barrier by building a data engine powered by 2D foundation models. Our data engine automatically annotates any number of object parts: 1755$\times$ more unique part types than existing datasets combined. By training on our annotated data with a simple contrastive objective, we obtain an open-world model that generalizes to any part in any object based on any text query. Even when evaluated zero-shot, we outperform existing methods on the datasets they train on. We achieve $\mathbf{260\%}$ improvement in mIoU and boost speed by 6$\times$ to 300$\times$. Our scaling analysis confirms that this generalization stems from the data scale, which underscores the impact of our data engine. Finally, to advance general-category open-world 3D part segmentation, we release a benchmark covering a wide range of objects and parts. Project website: https://ziqi-ma.github.io/find3dsite/
Figure 1: Find3D is the first general-category 3D model that can segment any part of any object with any text query. We achieve this by building a scalable Data Engine powered by 2D foundation models – SAM & Gemini – that automatically annotates 3D assets from the web. Using the labeled data, Find3D trains a transformer-based point cloud model with a contrastive training recipe. Our method works on diverse 3D objects and parts, e.g. the easel, the imaginary animal, and the ceiling fan.
@@ 1 Introduction
Is it possible to build foundation models in 3D? For text and image modalities, we have seen that strong, general models come from internet-scale training data. In the absence of such large-scale 3D datasets, can we attempt to replicate this success in 3D?
In this paper, we provide an answer to this question. We show that when you tackle the data challenge, you can get a strong model with a simple, general training recipe. This approach not only unlocks generalization to unseen objects with a $\mathbf{260\%}$ improvement in mIoU, but even outperforms prior methods on the datasets they train on.
Prior works, heavy on dataset-specific customization, suffer from poor generalization because existing datasets are small and homogeneous. For example, ShapeNet-Part [27] contains only 16 categories and makes assumptions such as “all chairs face right”. Evaluation on such limited datasets implicitly encourages dataset-specific customization, which is not the path towards generalization. Moreover, many prior works use pipelines that are computationally expensive and cannot scale to larger training sets.
In this paper, our goal is to: 1) establish a scalable data engine that can generate useful labels for any number of 3D assets; and 2) show that having a large-scale training set enables strong generalization with a simple training recipe, without any customizations for specific datasets such as per-category prompt and viewpoint search [33], category-specific finetuning [32, 21], multi-pass inference customization with predefined part ranking logic [32], or slow test-time pipelines [20]. Conceptually, our findings mirror those in other domains such as text, where using general training recipes at scale leads to powerful and general foundation models.
Concretely, as shown in Fig. 1, we enable scaling in 3D by building a data engine that automatically annotates synthetic 3D assets on the internet, yielding $\mathbf{2.1}$ million part annotations of 761 object categories. Our dataset contains $124615$ unique part types, which is over $\mathbf{1775\times}$ the number of unique part types in existing datasets combined (ShapeNet-Part [27] and PartNet-E [12] contain 71 unique part types combined). To leverage such large-scale data, we devise a contrastive training objective to handle part hierarchy and ambiguity. Our model takes in a point cloud and predicts a queryable semantic feature for every point. The features are in the latent embedding space of a CLIP-like [15] model, so that they can be queried with any free-form text by calculating pointwise cosine similarities with the query embedding.
This approach yields a model that can segment any part of any object, with any text query. We highlight the following contributions:
@@
  • We build the first model for 3D segmentation that is simultaneously open-world, cross-category, part-level and feed-forward. We achieve $\mathbf{260\%}$ improvement in mIoU and 6$\times$ to over 300$\times$ the inference speed compared to existing methods.
  • We release a benchmark for evaluating open-world 3D part segmentation for diverse objects, with 5$\times$ more unique part types than the largest existing benchmark.
Figure 2: The Data Engine. (a) We render Objaverse assets into multiple views and pass each rendering to SAM with gridpoint prompts for segmentation. For each mask, we query Gemini for the corresponding part name, which gives us (mask, text) pairs. We embed the part name into the latent embedding space of a vision and language foundation model such as SigLIP. We back-project mask pixels to obtain the points associated with each label embedding, yielding (points, text embedding) pairs. (b) Example annotations by the data engine.
@@ 2 Related Work
Closed-world 3D segmentation. 3D segmentation has been studied primarily in a closed world and with a coarse granularity that cannot go below whole objects. In specific settings such as indoor scenes or self-driving, state-of-the-art models are starting to achieve better generalization by training on multiple datasets, such as Mask3D [18] and the PointTransformer series [30, 23, 22]. However, these models are still domain-specific, and can only segment whole objects rather than parts. Part-level segmentation is less studied. Early efforts started with the ShapeNet-Part dataset [27] (16 object classes, $\leq 6$ parts per object). PartNet-E introduces articulated objects but is still limited to only 45 categories. Due to the limited number of categories and shared orientations (e.g., chairs all facing right), state-of-the-art part-level models [14, 9] cannot generalize well. Our work tackles both the challenges of generalization and granularity – our model is part-level, and can segment any object part in an open-world setting.
3D aggregation methods based on 2D renderings. With the progress of vision language models in 2D image understanding, some works directly assemble these models to obtain an “aggregated” 3D understanding without training a 3D model. An exemplary aggregation method uses multiview renderings of 3D scenes or objects, obtains their features in 2D based on models like CLIP [15], SAM [6], or GLIP [7], and combines them in 3D based on projection geometry. On the whole object level, such methods include OpenMask3D [20]. On the part level, such methods include PointCLIP [29], PointCLIPV2 [33], PartSLIP [8] and PartSLIP++ [32] for point clouds, and SATR [1] for meshes. These models lack 3D geometry information and suffer from inconsistency across views. Furthermore, these methods are slow because they perform many inferences and the aggregation logic at test time. Our method, which predicts in 3D with a single inference, is significantly faster. Our method also achieves stronger performance and better robustness to pose changes by leveraging 3D geometry information.
Test-time optimization. Test-time optimization methods combine features from 2D models with a 3D representation, such as NeRF or Gaussian Splatting. At test time, these methods optimize the 3D representation with the 2D-sourced features attached. LERF [3], Distilled Feature Field [19], and Garfield [4] are based on radiance fields. Feature3DGS [31] is based on Gaussian splatting. These methods need to be optimized per scene (or per object), which can be slow (several minutes). Moreoever, their part-level capabilities have not been well-studied. Our method, feed-forward in nature, provides much faster inference with better performance.
@@ 3 Method
Figure 3: Find3D: an open-world part segmentation model. Find3D takes in a point cloud, voxelizes and serializes the points via space-filling curves into a sequence. The sequence is passed through a transformer architecture which returns a pointwise feature that is in the embedding space of a vision and language foundation model, denoted by $\mathbb{T}$. These features can be queried with any free-form text. Find3D is trained with a contrastive objective. For each (points, text embedding) label from the data engine, we use the averaged feature of these points as the predicted embedding, and pair it with the text embedding to form a positive pair in the contrastive loss.
@@
Formally, given a point cloud $C=\{\mathbf{p_{1}},...{\mathbf{p_{n}}}\}$ with color and normals, for any point $\mathbf{p_{i}}=(x,y,z,n_{x},n_{y},n_{z},r,g,b)\in\mathbb{R}^{9}$, we want to find a semantic feature $f_{i}\in\mathbb{R}^{d}$ which belongs in the same latent embedding space as a CLIP-like model, e.g., SigLIP. At inference time, for any text $s$, we can get its SigLIP embedding $T(s)$ and compute its cosine similarity with $f_{i}$, $cos(T(s),f_{i})$. For segmentation, Find3D assigns each point to the text query with the highest cosine similarity, and assigns “no label” if all queries yield negative similarity scores.
@@ 3.1 Data Engine
Obtaining large-scale 3D annotations for generic object categories with human-in-the-loop pipelines is onerous. We develop a scalable data engine that leverages annotations from 2D foundation models and geometrically unprojects them to 3D.
As illustrated in  Fig. 2, our data engine leverages SAM [6] and Gemini [17] to annotate 3D assets from Objaverse [2]. Since Objaverse assets do not have a fixed orientation, and Gemini provides higher-quality labels to objects seen in familiar orientations, we first prompt Gemini to select the best orientation based on 10 renderings (from different camera angles) of an object in each orientation. For the chosen orientation, we pass all renderings to SAM with grid point prompts. We discard masks that are too small (less than 350 pixels out of a 500$\times$500 image), too large (greater than 20% of all pixels), or with low confidence from SAM. We overlay each mask on the original image and ask Gemini to name the shaded part. Prompts are detailed in the appendix. Masks with the same label are merged. This process generates labeled (mask, text) pairs. We map each mask to a set of points in the point cloud based on projection geometry. To make the point features queryable by language, we align point features to the language embedding space of a pretrained model, such as SigLIP. We embed the label texts and use the text features as supervision.
The data engine processes 36044 Objaverse objects under LVIS categories selected by  [11, 10]. Each part can be annotated differently from different views, denoting various aspects of part, such as location (e.g., “bottom”), material (e.g., “snowball”), and function (e.g., “body”). Labels also have different levels of granularity. For example, in Fig. 2, one granularity is individual snowballs, and another granularity is the whole snowman. The diversity of our labels helps the model handle the inherent ambiguity in segmentation. Our data engine annotates 30K objects from 761 unique categories with 2.1 million parts in total. Our annotations contain 124615 unique part types, which is over $\mathbf{1775\times}$ the number of unique part types in existing datasets combined (ShapeNet-Part and PartNet-E contain 71 unique part types in total). Fig. 2 panel b shows some example annotations covering a wide range of part types and object geometries. We provide more annotation examples by our data engine in the appendix.
@@ 3.2 Open-World 3D Part Model
Architecture. Find3D adopts the PT3 [22] architecture that treats point clouds as sequences, as illustrated in Fig. 3. To align the point features into the latent embedding space of SigLIP, we append a lightweight 4-layer MLP to the last layer of the transformer. This returns a 768-dimension feature per point. Our model contains 46.2 million parameters.
@@
Training. Leveraging the diverse annotations from our data engine requires some care. We cannot define a direct pointwise loss because: 1) The same point can have multiple labels that denote various aspects of a part such as location, material, and function. Some labels may also be incorrect; and 2) Many points are unlabeled - as shown in Fig. 3 (right), each mask only labels points visible from one camera view, and thus parts are likely to be labeled partially.
@@
The challenge of partial labels can be resolved if the model can map features based on 3D geometry: points on the same ball should share similar features, and if we align the features of some points on that ball correctly to the text embedding, the other points’ features should also be aligned. The challenge of multiple labels can be resolved by the contrastive formulation: each point’s feature is encouraged to be close to the embeddings of all its labels, which allows for flexible text queries at inference time. As illustrated in  Fig. 3 (right), we define the contrastive pairing as follows: for each label, the ground truth is the SigLIP embedding of the text. The predicted value is the average feature of all points that correspond to the label. This pooling can also be regarded as a way to “denoise” the labels – while an individual point might be affected by conflicting or incorrect labels, it is unlikely that all points are subjected to the same error.
@@
Formally, our data engine provides (points, text embedding) labels, which we denote as $(C_{i},T(\text{label}_{i}))$ where $C_{i}$ is a subset of the point cloud that this label corresponds to, and $T(\text{label}_{i})$ is the label embedding. We denote the pooled feature from the labeled points as $f(C_{i})$, where $f$ is our model. We define the contrastive loss as follows:
$l_{i}=-\log\frac{\exp(f(C_{i})\cdot T(\text{label}_{i}))}{\sum_{j=1}^{|\mathcal{B}|}\exp(f(C_{i})\cdot T(\text{label}_{j}))}$ (1)
@@
where $\mathcal{B}$ denotes all labels of all objects in a batch. For training, we use a batch size of 64 objects, corresponding to $\sim 3000$ positive pairs per batch.
@@
To achieve generalization, in addition to training on diverse data provided by the data engine, we also apply data augmentations, including random rotation (implemented as sequential random rotation along all three axes), scaling, flipping, jittering, chromatic auto contrast, chromatic translation, and chromatic jitter. These augmentations help avoid over-reliance on object poses and color, and nudges the model to take up 3D geometric cues. We perform a 90:10 train-validation split on the 27552 objects provided by the data engine, and train with the Adam optimizer [5] with a cosine annealing learning rate schedule, starting at 0.0003 and ending at 0.00005 over 80 epochs.